OpenStudy (tennisthegreat123):

The perimeter of a rectangle is 68 feet. The ratio of its length to it's width is 9:8 What are the dimensions of the rectangle? a. 20 ft by 14 ft b. 18 ft by 16 ft c. 24 ft by 10 ft d. 22 ft by 12 ft

OpenStudy (tennisthegreat123):

@563blackghost

563blackghost (563blackghost):

What was the equation we had came up with in the previous question?

OpenStudy (tennisthegreat123):

2 (9x) + 2 (8x) = 68

563blackghost (563blackghost):

Correct ^^ Now we need to simplify this so we can find x in the equation so....we would first simplify \(\Large{2(9x)}\) and \(\Large{2(8x)}\) in the equation.... \(\Huge{?x+?x=68}\) What would \(\Large{2(9x)}\) and \(\Large{2(8x)}\) equal?

OpenStudy (tennisthegreat123):

18x + 16x = 68

563blackghost (563blackghost):

Correct ^^ Now we need to add like terms....so we would add together \(\Large{18x}\) and \(\Large{16x}\).... \(\Huge{?x=68}\)

OpenStudy (tennisthegreat123):

34x = 68

563blackghost (563blackghost):

Once again correct xD Now we would need to divide to get x by itself and find out what x equals...so we would divide by \(\Large{34}\) and what we do to one side we do to the other... \(\Huge{\frac{34x}{34}=\frac{68}{34} \rightarrow x=?}\) \(\Large{\frac{34}{34}}\) Would cancel out and x would be left so what would \(\LARGE{\frac{68}{34}}\) equal?

OpenStudy (tennisthegreat123):

x=2

563blackghost (563blackghost):

Correct ^^ So now that we know that x equals 2 we would input 2 into the equation and find out what the width would equal and the length.... \(\Huge{8(2)=w}\) \(\Huge{9(2)=l}\)

OpenStudy (tennisthegreat123):

16= w 18= l

563blackghost (563blackghost):

Correct ^^ So what would be the answer?

OpenStudy (tennisthegreat123):

B :)

563blackghost (563blackghost):

Great Job ^^

OpenStudy (tennisthegreat123):

ty for ur help once again

563blackghost (563blackghost):

np ^^ im glad to be of help...if you ever need help in the future just tag me in your question